OpenStudy (anonymous):

Given the equation x^2+y^2+20x-10y-12=0 , (a) find the center and radius; (b) graph; (c) find the intercepts, if any. Please show all of your work.

OpenStudy (anonymous):

complete the square for this one

OpenStudy (anonymous):

\[x^2+y^2+20x-10y-12=0\] \[x^2+20x+y^2-10y=12\] \[(x+10)^2+(y-5)^2=12+10^2+5^2=137\] so now it i looks like \[(x-h)^2+(y-k)^2=r^2\] where the center is \[(h,k) \text{ or } (-10,5)\] and the radius is \[\sqrt{137}\]

OpenStudy (anonymous):

ty